OpenStudy (anonymous):

Simplify this expression. (x6)3 = A. x2 B. x3 C. x9 D. x18

OpenStudy (imtiaz7):

it is (x^6)^3?

OpenStudy (anonymous):

...What's 6 times 3?

OpenStudy (anonymous):

Oh wait, it might be an exponent.

OpenStudy (umarsback):

x9*

OpenStudy (anonymous):

yea its an exponet

OpenStudy (anonymous):

I think its x^18?

OpenStudy (anonymous):

\[(x ^{6})^{3?}\]

OpenStudy (umarsback):

haha did she take my answer seriously

OpenStudy (umarsback):

6*3 = ?

OpenStudy (anonymous):

@amishjeb yea its that.

OpenStudy (imtiaz7):

x^18

OpenStudy (anonymous):

\[x ^{18}\]

OpenStudy (anonymous):

\[x^{2} * x^{3} = x^{5}\] if its raised to the power like yours you multiply

OpenStudy (imtiaz7):

(x6)3 = x^6*3 = x^18